OpenStudy (anonymous):

The length of a rectangle is one less than twice its width. If its perimeter is 34", what are the dimensions?

OpenStudy (texaschic101):

P = 2L + 2W P = 34 L = 2W - 1 now we sub 34 = 2(2W - 1) + 2W 34 = 4W - 2 + 2W 34 + 2 = 4W + 2W 36 = 6W 6 = W now sub 6 in for W in either equation to find L L = 2W - 1 L = 2(6) - 1 L = 12 - 1 L = 11 check... P = 2W + 2L 34 = 2(6) + 2(11) 34 = 12 + 22 34 = 34 (correct) solution : W = 6 and L = 11

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

No I understand. Thank you

OpenStudy (texaschic101):

YW :)